The number of non-text books read by college students ranges from 7 to 49. Using B as the variable, write an absolute value inequality that corresponds to this range.
Answer by stanbon(75887)

The number of non-text books read by college students ranges from 7 to 49. Using B as the variable, write an absolute value inequality that corresponds to this range.
---
Find the center of the range: (7+49)/2 = 28
Find the distance from the center to the end point: 49-28 = 21
----
Absolute inequality:
|B-28| <= 21
====================
Cheers,
Stan H.
===========